# SAM Decoding: Speculative Decoding via Suffix Automaton

### v2

@@ Appendix C Additional Ablation Experiments
Figure 5: the percentage of usage times and mean accept tokens of different draft modules.
@@
The inference process of SAM-Decoding is divided into five stages: prefill, draft generation, decoding, verification, and updating. During the prefill stage, the model processes the input prompt to establish an initial state. In the draft generation stage, a draft is produced based on this initial state. The decoding stage involves the model further processing this draft. Next comes verification, where the correct parts of the draft are evaluated based on the information processed during the decoding stage. Finally, the update phase modifies the state of the model based on the valid parts of the draft. Figure 4 illustrates the proportion of time each stage consumes within the SAM-Decoding[T] process base on Spec-Bench. As shown, the decoding stage takes up the largest portion of time, accounting for 65.4% of the entire process. This is followed by the verification stage, which occupies 23.4% of the total time. The updating stage requires 6.3% of the time, whereas the draft generation stage contributes only 0.6% to the overall duration. Additionally, the prefill stage comprises 4.2% of the total processing time.
Figure 5 shows the usage times of different draft modules of SAM-Decoding[T] on Spec-Bench and the corresponding average draft accept length. It can be seen that in 85.96% of the cases, due to insufficient matching length, we generate drafts based on the auxiliary method, corresponding to an average accept length of 2.51, while in the remaining 11.59% and 2.45% of the cases, the dynamic suffix automaton and static suffix automaton are used to generate drafts, corresponding to average accept lengths of 6.57 and 3.39, respectively.

### v3

@@ Appendix C Additional Ablation Experiments
Figure 9: the percentage of usage and mean accept tokens of different draft modules.
@@
The inference process of SAM-Decoding is divided into five stages: prefill, draft generation, decoding, verification, and updating. During the prefill stage, the model processes the input prompt to establish an initial state. In the first draft generation stage, a draft is produced based on this initial state. The decoding stage involves the model further processing this draft. Next comes verification, where the correct parts of the draft are evaluated based on the information processed during the decoding stage. Finally, the update phase modifies the state of the model based on the valid parts of the draft. Figure 8 illustrates the proportion of time each stage consumes within the SAM-Decoding[T] process based on Spec-Bench. As shown, the decoding stage takes up the largest portion of time, accounting for 65.4% of the entire process. This is followed by the verification stage, which occupies 23.4% of the total time. The updating stage requires 6.3% of the time, whereas the draft generation stage contributes only 0.6% to the overall duration. Additionally, the prefill stage comprises 4.2% of the total processing time.
Figure 9 shows the usage frequency of different draft modules of SAM-Decoding[T] on Spec-Bench and the corresponding average draft accept length. It can be seen that in 85.96% of the cases, due to insufficient matching length, we generate drafts based on the auxiliary method, corresponding to an average accept length of 2.51, while in the remaining 11.59% and 2.45% of the cases, the dynamic suffix automaton and static suffix automaton are used to generate drafts, corresponding to average accept lengths of 6.57 and 3.39, respectively.
Finally, Table 9 shows the inference speed of different methods based on Vicuna-7B-v1.3 on NVIDIA A800 GPU. It can be seen that SAM-Decoding can still effectively combine Token Recycling and EAGLE-2 to achieve higher inference speed, which shows the effectiveness of our approach for different devices.